What is the smallest composite number over 10,000 whose digits sum to the same value as the sum of the digits of its prime factorization?

Found 2 solutions by Alan3354, RAY100:
Answer by Alan3354(69443)

10201 = 101*101
Might be it.

Answer by RAY100(1637)

Try this, assume that the digits are the same,
.
xx,xxx,,,,,or x+x+x+x+x (Sum of Digits)
.
if y = factor,,,,use 5 factors(somewhat arbitrary),again all factors same
,
As a 1st est,,,5th root of 10000 = 6.3,, therefore use 7
,
7^5 = 16,807,,,1+6+8+0+7 = 22 vs 5*7 =35
,
trying 8, 8^5 = 32768,,,,3+2+7+6+8=26, vs 5*8=40
.
from this pattern we see that the power should increase, and the base degrease
,
following this pattern, bring us to 3^9= 19683
Sum of Digits = 1+9+6+8+3 = 27,,,and 9*3 =27
.
Might be a number closer to 10,000, but this works